4.4 Gestational diabetes mellitus (GDM) is a condition that causes serious adverse perinatal outcomes in pregnancy. It is estimated that approximately 11.7% of women have GDM during pregnancy. A clinic trial was initiated to test a screening method for GDM among 37 pregnant women who had GDM and 33 pregnant women without GDM. The results are listed below Test (+) 36 Test (-) Total 37 Women with GDM Women without GDM 31 a) What is the sensitivity? b) What is the specificity? c) What is the probability of false negative? d) What is the probability of false positive? e) What is the prior probability of GDM? f) Suppose a pregnant woman has not yet seen a gynecologist/obstetrician but she has a positive GDM screening test. What is the probability this pregnant woman has GDM?
g) If the woman in f) had a negative test result, how likely is that she does not have G
Answer #1
+ve -ve total
GDM 36 1 37
no GDM 2 31 33
total 38 32 70

a) sensitivity is true positve rate = P(+ve | GDM)=36/37=               0.9730
              
b) specificity is true negative rate = P(-ve|no GDM) = 31/33=               0.9394
              
c) P(false negative) = 1/70=               0.0143
              
d) P(false positive) = 2/70=               0.0286
e) prior probability of GDM = 37/70=   0.5286
  
f) P(GDM| +ve test) = P(GDM and +ve test)/P(+ve test) = 36/38=   0.9474
  
g) P(no GDM|-ve test)=P(no GDM and -ve test)/P(no GDM) = 31/32=   0.9688